# Meson dynamics from locally exciting a particle-conserving $Z_{2}$ lattice gauge theory

### v1

@@ IV Effect of confinement field on meson dynamics
When $h/J\gg 1$, the relative coordinate is restricted to values around $r=1$. The mean value of the relative coordinate oscillates as a function of time. In Appendix A.2, we show $r_{\text{avg}}=1+J^{2}/(2h^{2})+J^{2}\sin^{2}(ht)/(2h^{2})$. We again get frequency of oscillations, $\omega=2h$ but the spread of average meson size around its initial value ($r=1$) scales as $(h/J)^{-2}$. We note that the $(h/J)^{-1}$ dependence of $r^{\prime}_{\text{avg}}$ (seen in Fig. 4) only captures an overall trend for moderate to large values of $h/J$. For very large values of $h/J$, the curve is expected to follow an $(h/J)^{-2}$ dependence.
@@
Figure 4: Long-time average size of propagating mesons, $r^{\prime}_{\text{avg}}$, vs confinement field, $h/J$ shown by open markers. The solid blue line is a fit to a function, $r^{\prime}_{\text{avg}}-1\propto 1/h$. Frequency of oscillations, $\omega$ of average meson size, $r_{\text{avg}}$ vs field $h/J$ shown by solid black dots. The solid red line is a fit of $\omega$ as a linear function of $h$.
@@ V Effect of the initial state’s energy on meson dynamics
Figure 5: A central excitation in a superposition of an $r=1$ and $r=2$ meson, tuned by the angle $\theta$ of one of the gauge spins. The gray circles show a finite probability between 00 and $1$ of having a particle controlled by $\theta$. The excitation energy is, $E(\theta)=\sin\theta+2h\sin^{2}(\theta/2)+2h$.
@@
Using the Hamiltonian in Eq. 3, the energy of this initial state is, $E(\theta)=\sin\theta+2h\sin^{2}(\theta/2)+2h$. The term $2h\sin^{2}(\theta/2)$ comes from the confining electric field energy of the $r=2$ meson amplitude and the $\sin\theta$ term denotes the kinetic energy from the hopping term of the Hamiltonian. We tune $\theta$ from 00 to $\pi$, varying the initial state from an $r=1$ meson continuously to an $r=2$ meson. The energy, $E(\theta)$, varies non-monotonically with $\theta$ as shown in Fig. 6. This is due to the kinetic energy contribution. Starting from $|\psi_{\theta}(0)\rangle$ at a fixed confining field, $h=1.1J$, we calculate the average size, $r^{\prime}_{\text{avg}}$, of propagating mesons at long times and plot it in Fig. 6 as a function of $\theta$. We find that $r^{\prime}_{\text{avg}}$ closely tracks the initial state’s energy, having the same functional dependence on $\theta$ as energy does. If the initial energy increases, $r^{\prime}_{\text{avg}}$ rises, indicating larger mesons. Note that although the initial state is spatially largest when $\theta=\pi$ (an $r=2$ meson), the average size of mesons is lower compared to the case when $\theta=3\pi/4$ where the initial energy is higher. This occurs because the initial kinetic energy also gets converted into the size of mesons.
Figure 6: At $h=1.1J$, initial state’s energy, $E$ (red dashed line) and average meson size, $r^{\prime}_{\text{avg}}$ (black dots are data and blue solid line is a fit) as a function of $\theta$. The meson size and energy have the same functional dependence on $\theta$. Higher energy leads to longer meson excitations that propagate.
Figure 7: Particle density on each site (x-axis) as a function of time, $Jt$ (y-axis), for $h=1.1J$ for different initial energies, $E$, and average meson size, $r^{\prime}_{\text{avg}}$. (a) $E=2.67J$ and $r^{\prime}_{\text{avg}}=1.6$, (b) $E=3.8J$ and $r^{\prime}_{\text{avg}}=1.98$ and (c) $E=4.78J$ and $r^{\prime}_{\text{avg}}=2.34$. We can see that as the average meson size, $r^{\prime}_{\text{avg}}$, increases with energy, the particle density spread becomes narrower. Longer mesons are propagating at slower speeds away from the center.
@@
Figure 8: The speed of propagating mesons, $v$, as a function of average meson size, $r^{\prime}_{\text{avg}}$, at (a) $h=1.1J$ and (b) $h=0.3J$. The meson size at a fixed $h$ is varied by varying the initial energy of the state. The speed $v$ and $r^{\prime}_{\text{avg}}$ are inversely related, making longer mesons slower. This suggests that longer mesons are effectively heavier due to the $Z_{2}$ gauge constraint. This is not only true when $h>J$ but persists for values where $h<J$.
@@
The speed of the propagating mesons is strongly affected by their size at a fixed confining field, $h/J$. We show the particle density evolution as a function of time for different initial energies and correspondingly different average meson sizes in Fig. 7. We can clearly see that as the average meson size increases, the particle density spread becomes narrower, indicating that longer mesons are propagating at slower speeds. In Fig. 8, we plot the speed of the propagating meson excitations as a function of their size, $r^{\prime}_{\text{avg}}$, at $h=1.1J$ and $h=0.3J$ by changing the initial state energy as discussed above. The speed is calculated by finding the slope, $v=dc_{s}/d(Jt)$ at long times from the $c_{s}$ vs $Jt$ graph (example shown in Fig. 3(b)). We find that meson size and speed are inversely related in both strong confinement ($h>J$) and weak confinement ($h<J$) cases. Longer mesons propagate more slowly. Thus a detector placed far away from the center would detect smaller mesons first and would have to wait for a longer time to detect larger mesons. This dependence of meson speed on meson size is due to the kinetic constraints arising from the linear confining potential. Following arguments similar to those in  [42, 53] we show how linear confinement leads to faster speeds for shorter mesons in the $h/J\gg 1$ limit. Furthermore we argue how this phenomenology remains valid in our scenario even when $h<J$.
@@
We first show how an $r=1$ meson would move faster than an $r=2$ meson and so on for longer mesons when $h>J$. Consider an $r=1$ meson at some large value of $h/J$. The entire meson can hop to its right only by a second order hopping process as shown in Fig. 9(a). This process changes its center-of-mass position by 1 but leaves the relative coordinate unchanged. By second order perturbation theory, the matrix element for this process is given by $-J^{2}/2h$. As the confining field $h/J$ increases, the $r=1$ mesons propagate slower.
@@
Now let’s consider an $r=2$ meson. In a similar vein, we can consider processes which lead to the meson hopping to its right. There are two possibilities at second order. The first is shown in Fig. 9(b) where the right particle hops to the right first and then the left particle follows. The matrix element for this process is $-J^{2}/2h$ similar to the $r=1$ meson. The second possibility shown in Fig. 9(c) is similar, but here the left particle hops first, followed by the right particle. The matrix element of this process is $J^{2}/2h$. The matrix elements of these two processes are equal and opposite in sign. They cancel out and thus the $r=2$ meson cannot hop by a second order hopping process. The leading order process that can cause motion is fourth order, shown in Fig. 9(d). Here the right particle hops two sites and then the left particle follows. The center of mass changes by 2 and the matrix element for this process is $-J^{4}/16h^{3}$. The propagation speed for an $r=2$ meson would be $v\sim J^{4}/8h^{3}$ as opposed to $J^{2}/2h$ for an $r=1$ meson. At large values of $h/J$, the $r=2$ meson would be much slower than the $r=1$ meson. Using a similar argument, one can show that an $r=3$ meson can only propagate by a sixth order process and is even slower than $r=2$ mesons.
[…]
@@ VI Proposed experimental implementation using a spin model
$H=-\sum_{i}\frac{(1-\sigma^{z}_{i-1,i}\sigma^{z}_{i+1,i+2})}{2}\sigma^{x}_{i,i+1}+h\sum_{i}\sigma^{z}_{i,i+1}$ (11)
In the particle vacuum, all the spins point down. A particle excitation corresponds to a domain wall in the spins. The first term in the spin model encodes particle hopping from an occupied site (presence of a spin domain wall) to an empty site (absence of a domain wall) by flipping the link spin in the path. The initial state in the dynamics considered here are simply product states of the spins. The initial spin configuration for the $r=1$ meson initial state is shown in Fig. 1(b). When we tune the initial state energy by introducing a finite amplitude of $r=2$ meson, we simply need to prepare the spin configuration shown in Fig. 5. Here the central spin is pointing up and the spin next to it is rotated along the $y$-axis by an angle $\theta$ using a single qubit $y$-rotation.
For studying dynamics, the spin model in Eq. 11 can ber Trotterized on a quantum computer. The only non-traditional part of this model is the three-spin term encoding particle hopping. It requires a native three qubit gate or its decomposition into a set of two-qubit gates. All the quantities discussed in our work can be accessed by analyzing the snapshots obtained by measuring all the spins at any time step. A particular snapshot would contain two spin domain walls, indicating the two particles. The distance between the domain walls averaged over snapshots gives the average meson size, $r_{\text{avg}}$. The center-of-mass distribution’s width, $c_{s}$ can also be found by recording the position of domain walls that occur at any time step in the system.
@@ VII Summary and Outlook
We performed exact diagonalization numerics to simulate the dynamics and found that the long-time state is characterized by the propagation of a superposition of different sizes of mesons. The mesons show size oscillations in time. The average meson size depends inversely on the confinement field while the frequency of oscillations is proportional to the confining field. The meson size oscillations become larger as the confinement field strength decreases. These size oscillations are similar to the particle tracks seen after wavepacket collisions in lattice gauge theories [55, 51]. We were able to explain these behaviors analytically by solving an associated Wannier-Stark problem in the large field ($h/J$) limit.
@@
We calculated the speed of the meson propagation as a function of their size and found that longer mesons are effectively heavier and propagate more slowly. Longer mesons carry higher gauge field energy excitations giving them a larger effective mass. Due to different propagation speeds, mesons of different sizes spatially get separated as we evolve in time. We found that these results hold not only when $h/J$ is large but also for smaller values of $h/J$. We showed that for any finite linear confining potential, the leading order meson hopping matrix elements show exponential decay for mesons beyond a certain length. This threshold length rises as the confinement field becomes weaker and continues all the way until we approach a continuum limit. Throughout the study, we only consider simple initial states that create a localized high energy excitation region, making these phenomena experimentally accessible in the near-term quantum simulators.
@@
It would be interesting to explore whether these features occur in generic linear confining gauge theories in an arbitrary dimension or if they are just a special case of one dimensional gauge theories. An interesting direction is to study higher dimensional analogs of the problem we studied and the effects of introducing particle non-conserving terms. We can expect to see richer phenomenology even in this simple set-up as longer mesons can either break up into shorter mesons or create other kinds of gauge field excitations such as magnetic field like excitations.

### v2

@@ IV Effect of confinement field on meson dynamics
When $h/J\gg 1$, the relative coordinate is restricted to values around $r=1$. We note that in this limit, the initial state is nearly an eigenstate of the hamiltonian and remains stationary for long times. However there would still be small fluctuations in the mean value of the relative coordinate as a function of time. In Appendix A.2, we show $r_{\text{avg}}=1+J^{2}/(2h^{2})+J^{2}\sin^{2}(ht)/(2h^{2})$. We again get frequency of oscillations, $\omega=2h$ but the spread of average meson size around its initial value ($r=1$) scales as $(h/J)^{-2}$. For very large values of $h/J$, the curve is expected to follow an $(h/J)^{-2}$ dependence.
@@
We note that the $(h/J)^{-1}$ dependence of $r^{\prime}_{\text{avg}}$ (seen in Fig. 4) only captures an overall trend for moderate values of $h/J$. This scaling connects the analytically expected scalings in the $h/J\ll 1$ and $h/J\gg 1$ limits. Our numerical simulations focus on these moderate $h/J$ values as this is an interesting non-trivial regime where the simple analytical approximations presented earlier break down. In Appendix B, we show a log-log plot of $(r^{\prime}_{\text{avg}}-1)$ vs $h/J$ where we show how the scaling changes from $(h/J)^{-1}$ to $(h/J)^{-2}$ as we move towards much larger values of $h/J$.
Figure 4: Long-time average size of the propagating meson, $r^{\prime}_{\text{avg}}$, vs confinement field, $h/J$ shown by open markers. The solid blue line is a fit to a function, $r^{\prime}_{\text{avg}}-1\propto 1/h$. The error bars show the long time average of the standard deviation of the meson size, $r^{\prime}_{\text{sd}}$. It shows the range of meson size fluctuations. Frequency of oscillations, $\omega$ of average meson size, $r_{\text{avg}}$ vs field $h/J$ shown by solid green triangles. The solid red line is a fit of $\omega$ as a linear function of $h$.
@@ V Effect of the initial state’s energy on meson dynamics
Figure 5: A central excitation in a superposition of an $r=1$ and $r=2$ meson, tuned by the angle $\theta$ of one of the gauge spins. The gray circles show a finite probability between 0 and 11 of having a particle controlled by $\theta$. The excitation energy is, $E(\theta)=\sin\theta+2h\sin^{2}(\theta/2)+2h$.
@@
Using the Hamiltonian in Eq. 3, the energy of this initial state is, $E(\theta)=J\sin\theta+2h\sin^{2}(\theta/2)+2h$. The term $2h\sin^{2}(\theta/2)$ comes from the confining electric field energy of the $r=2$ meson amplitude and the $\sin\theta$ term denotes the kinetic energy from the hopping term of the Hamiltonian. We tune $\theta$ from 0 to $\pi$, varying the initial state from an $r=1$ meson continuously to an $r=2$ meson. The energy, $E(\theta)$, varies non-monotonically with $\theta$ as shown in Fig. 6(a),(b) for $h/J=1.1,0.3$. This is due to the kinetic energy contribution. Starting from $|\psi_{\theta}(0)\rangle$ at a fixed confining field, we calculate the average size, $r^{\prime}_{\text{avg}}$, of the propagating meson at long times and plot it in Fig. 6(a),(b) as a function of $\theta$ for $h/J=1.1$ and $h/J=0.3$ respectively. We find that $r^{\prime}_{\text{avg}}$ closely tracks the initial state’s energy in both cases, having the same functional dependence on $\theta$ as energy does. If the initial energy increases, $r^{\prime}_{\text{avg}}$ rises, indicating a larger meson. Note that although the initial state is spatially largest when $\theta=\pi$ (an $r=2$ meson), the average size of mesons is lower compared to the case when $\theta=3\pi/4$ where the initial energy is higher. This occurs because the initial kinetic energy also gets converted into the size of mesons.
Figure 6: Initial state’s energy, $E$ (red dashed line) and average meson size, $r^{\prime}_{\text{avg}}$ (black dots are data and blue solid line is a fit) as a function of $\theta$ for (a) $h/J=1.1$ and (b) $h/J=0.3$. The meson size and energy have the same functional dependence on $\theta$ in both cases. Total energy has two contributions, kinetic energy from the hopping term and the electric energy from the gauge field confinement term. Higher total energy leads to longer meson excitation that propagates.
Figure 7: Particle density on each site (x-axis) as a function of time, JtJt (y-axis), for $h=1.1J$ for different initial energies, $E$, and average meson size, $r^{\prime}_{\text{avg}}$. (a) $E=2.67J$ and $r^{\prime}_{\text{avg}}=1.6$, (b) $E=3.8J$ and $r^{\prime}_{\text{avg}}=1.98$ and (c) $E=4.78J$ and $r^{\prime}_{\text{avg}}=2.34$. We can see that as the average meson size, $r^{\prime}_{\text{avg}}$, increases with energy, the particle density spread becomes narrower. Longer mesons are propagating at slower speeds away from the center.
@@
Figure 8: The speed of the propagating meson, $v$, as a function of average meson size, $r^{\prime}_{\text{avg}}$, at (a) $h=1.1J$ and (b) $h=0.3J$. The meson size at a fixed $h$ is varied by varying the initial energy of the state. Each data point corresponds to the corresponding data point for angle $\theta$ that varies the initial energy in Fig. 6. The speed $v$ and $r^{\prime}_{\text{avg}}$ are inversely related, making longer mesons slower. This suggests that longer mesons are effectively heavier due to the $Z_{2}$ gauge constraint. This trend persists for values where $h<J$.
@@
The speed of the propagating mesons is strongly affected by their size at a fixed confining field, $h/J$. We show the particle density evolution as a function of time for different initial energies and correspondingly different average meson sizes in Fig. 7. We can clearly see that as the average meson size increases, the particle density spread becomes narrower, indicating that longer mesons are propagating at slower speeds. In Fig. 8, we plot the speed of the propagating meson excitations as a function of their size, $r^{\prime}_{\text{avg}}$, at $h=1.1J$ and $h=0.3J$ by changing the initial state energy as discussed above. The speed is calculated by finding the slope, $v=dc_{s}/d(Jt)$ at long times from the $c_{s}$ vs $Jt$ graph (example shown in Fig. 3(b)). We find that meson size and speed are inversely related in both strong confinement ($h>J$) and weak confinement ($h<J$) cases. Longer mesons propagate more slowly. Thus a detector placed far away from the center would detect smaller mesons first and would have to wait for a longer time to detect larger mesons. This dependence of meson speed on meson size is due to the kinetic constraints arising from the linear confining potential. This phenomena is known to be nonperturbative in the ratio $h/J$ [35]. Following arguments similar to those in [42, 53] we show how linear confinement leads to faster speeds for shorter mesons in the $h/J\gg 1$ limit. Furthermore we use a matrix element based approach to quantitatively show how this phenomenology is manifest in the dynamics when $h<J$.
@@
We first show how an $r=1$ meson would move faster than an $r=2$ meson and so on for longer mesons when $h>J$. Consider an $r=1$ meson at some large value of $h/J$. The entire meson can hop to its right only by a second order hopping process as shown in Fig. 9(a). This process changes its center-of-mass position by 1 but leaves the relative coordinate unchanged. By second order perturbation theory, the matrix element for this process is given by $J^{2}/2h$. As the confining field $h/J$ increases, the $r=1$ mesons propagate slower.
@@
Now let’s consider an $r=2$ meson. In a similar vein, we can consider processes which lead to the meson hopping to its right. There are two possibilities at second order. The first is shown in Fig. 9(b) where the right particle hops to the right first and then the left particle follows. The matrix element for this process is $J^{2}/2h$ similar to the $r=1$ meson. The second possibility shown in Fig. 9(c) is similar, but here the left particle hops first, followed by the right particle. The matrix element of this process is $J^{2}/2h$. The matrix elements of these two processes are equal and opposite in sign. They cancel out and thus the $r=2$ meson cannot hop by a second order hopping process. The leading order process that can cause motion is fourth order, shown in Fig. 9(d). Here the right particle hops two sites and then the left particle follows. The center of mass changes by 2 and the matrix element for this process is $J^{4}/16h^{3}$. The propagation speed for an $r=2$ meson would be $v\sim J^{4}/8h^{3}$ as opposed to $J^{2}/2h$ for an $r=1$ meson. At large values of $h/J$, the $r=2$ meson would be much slower than the $r=1$ meson. Using a similar argument, one can show that an $r=3$ meson can only propagate by a sixth order process and is even slower than $r=2$ mesons.
Figure 9: (a) Second-order hopping process (path I followed by path II) of an $r=1$ meson with matrix element $J^{2}/(2h)$. (b),(c) Two possible second order hopping processes of an $r=2$ meson with path I followed by path II and vice versa. The matrix elements for the two processes are equal and opposite and hence cancel. (d) A fourth order hopping process (I $\rightarrow$ II $\rightarrow$ III $\rightarrow$ IV) with matrix element $J^{4}/16h^{3}$ that changes the center of mass position of an $r=2$ meson by 2.
@@ VI Proposed experimental implementation using a spin model
$H=-J\sum_{i}\frac{(1-\sigma^{z}_{i-1,i}\sigma^{z}_{i+1,i+2})}{2}\sigma^{x}_{i,i+1}+h\sum_{i}\sigma^{z}_{i,i+1}$ (11)
In the particle vacuum, all the spins point down. A particle excitation corresponds to a domain wall in the spins. The first term in the spin model encodes particle hopping from an occupied site (presence of a spin domain wall) to an empty site (absence of a domain wall) by flipping the link spin in the path. We can also obtain this model from the infinite Ising coupling limit of the 1D transverse field Ising model. The Ising model Hamiltonian is
$H=-J\sum_{i}\sigma^{x}_{i,i+1}+h\sum_{i}\sigma^{z}_{i,i+1}+m\sum_{i}\sigma^{z}_{i-1,i}\sigma^{z}_{i,i+1}$ (12)
If the coefficient of the Ising coupling term $m$ is infinitely large, the energy cost to change the number of domain walls is infinite. In dynamics, the number of domain walls (particle excitations) would remain conserved. The effective Hamiltonian in that limit would correspond to the Hamiltonian in  11. For large but finite $m$, although the number of domain walls are not exactly conserved, we expect our results to continue to hold in the Ising model based on previously shown prethermalization arguments [35, 60].
The initial state in the dynamics considered here are simply product states of the spins. Spin product states in the context of lattice gauge theories have been prepared in superconducting circuits [4, 7, 9] and trapped ions [12, 13]. The initial spin configuration for the $r=1$ meson initial state is shown in Fig. 1(b). When we tune the initial state energy by introducing a finite amplitude of $r=2$ meson, we simply need to prepare the spin configuration shown in Fig. 5. Here the central spin is pointing up and the spin next to it is rotated along the $y$-axis by an angle $\theta$ using a single qubit $y$-rotation gate.
For studying dynamics, the spin model in Eq. 11 can be Trotterized on a digital quantum computer. Alternatively, it can be studied on analog quantum simulators such as trapped ions by implementing the large-coupling limit of the Ising model in Eq. 12 [12, 13, 9]. In the context of digital simulations, the only non-traditional part of the spin model in Eq. 11 is the three-spin term encoding particle hopping. This comes from the gauge-matter coupling term in the lattice gauge theory model and it requires a native three qubit gate for unitary evolution. As this term is composed of Pauli operators, the three qubit term unitary can be decomposed into multiple two-qubit unitaries involving single qubit rotations and two-qubit gates involving at least one of the three qubits [61, 62]. We show this in Appendix C. Universal quantum computers can implement this term using their universal two-qubit gate set. In Appendix C, we discuss the fidelity of two-qubit gates needed as a function of system size, trotter step and evolution time to observe the phenomenology that we study.
All the quantities discussed in our work can be accessed by analyzing the snapshots obtained by measuring all the spins at any time step. A particular error-free snapshot would contain two spin domain walls, indicating the two particles. The distance between the domain walls averaged over snapshots gives the average meson size, $r_{\text{avg}}$. The center-of-mass distribution’s width, $c_{s}$ can also be found by recording the width of the distribution of domain wall positions that occur at any time step in the system. In Appendix C, we discuss the number of snapshots needed to observe the meson dynamics phenomena.
@@ VII Summary and Outlook
We performed exact diagonalization numerics to simulate the dynamics and found that the long-time state is characterized by the propagation of a meson that is a superposition of different meson sizes. The meson shows size oscillations in time. The average meson size depends inversely on the confinement field while the frequency of oscillations is proportional to the confining field. The meson size oscillations become larger as the confinement field strength decreases. These size oscillations are similar to the particle tracks seen after wavepacket collisions in lattice gauge theories [55, 51]. We were able to explain these behaviors analytically by solving an associated Wannier-Stark problem in the large field ($h/J$) limit.
@@
We calculated the speed of the meson propagation as a function of its size and found that a longer mesons is effectively heavier and propagates more slowly. Longer mesons carry higher gauge field energy excitations giving them a larger effective mass. Since the long time wavefunction of a single meson contains a superposition of different meson sizes, the correlation between meson size and speed causes a size dependent spatial separation of the superposition constituents. We found that these results hold not only when $h/J$ is large but also for smaller values of $h/J$. We showed that for any finite linear confining potential, the leading order meson hopping matrix elements show exponential decay for mesons beyond a certain length. This threshold length rises as the confinement field becomes weaker and continues all the way until we approach a continuum limit. Throughout the study, we only consider simple initial states that create a localized high energy excitation region, making these phenomena experimentally accessible in the near-term quantum simulators. Our initial states offer a clean setting that is less prone to state preparation errors.
@@
Our study is amenable to efficient classical simulations due to a restriction to the two particle case in a particle-conserving lattice gauge theory. However we must emphasize that this work’s primary goal is to understand some building blocks of high energy phenomena in lattice gauge theory theories in a simple, clean setting that does not require preparation of moving wavepackets. We were able to show that a localized high energy excitation region can give rise to interesting phenomenology of meson propagation and provide a controllable setting for a detailed characterization analysis.
Figuring out these building blocks is crucial as we begin to explore similar phenomenology in several closely related models that are hard to simulate classically. This includes models with particle non-conservation, higher dimensional settings and SU($N$) non-abelian lattice gauge theories. Studying the dynamics induced by localized high energy excitations in these scenarios would lead to an interesting interplay where the initial energy can be used to create gauge field excitations or additional matter particles such as mesons or even baryons in SU($N$) gauge theory models. Our work serves as a nascent yet important step in this direction where we can understand the dynamics of high energy phenomena in lattice gauge theories using only simple initial states as opposed to wavepacket collisions.